There are people walking their dogs at a park. there are 19 heads and 60 feet at the park.
a. write out the equation for this problem use P to stand for people and d to stand for dogs.
b. how many people are there at the park? how many dogs are there?

Answer by stanbon(75887)

There are people walking their dogs at a park. there are 19 heads and 60 feet at the park.
a. write out the equation for this problem use P to stand for people and d to stand for dogs.
Heads equation: p + d = 19
feet equation: 2p +4d = 60
----
Modify for elimination:
p + d = 19
p +2d = 30
----
Subtract and solve for "d":
d = 11 (# of dogs)
---
solve for "p":
p + d = 19
p + 11 = 19
p = 8 (# of people)
=========================
b. how many people are there at the park? 8
how many dogs are there? 11
=============================
cheers,
Stan H.
=====================